What volume of Cl2(g) measured as STP is produced when 7.65 g of HCl(aq) reacts?

1
Expert's answer
2022-05-02T04:53:03-0400

H2 + Cl2 = 2HCl


V(Cl2) = (m(HCl) * Vm(Cl2)) / 2Mr(HCl)

V(Cl2) = (7.65 g * 22.4 L/mol) / 2 * 36.46 g/mol = 2.35 L